# Visualization Tasks for Unlabelled Graphs

### v1

@@ 3.2 Action
Locate. Asks “where are the targets situated within the graph?”, and applies to the Subgraph and Constituent scopes. The location may involve proximity between targets of the same type, or between two different kinds of targets.
@@ 3.3 Targets
At the Multiple and Pair scopes, the target is simply the graph itself; that is, the core graph topology (Sec. 2) of a simple graph consisting of a single connected component. At the Pair scope there are two such targets, and at the Multiple scope there are many.
At the Single taxonomy scope, the target pertains to the entire graph, encompassing all of its nodes and edges. Many of the targets are names from the graph theory literature that describe global topological characteristics of a graph. A graph of $N$ nodes is a tree when it has $N-1$ edges and a complete graph when it has (${N\choose 2}$ edges. While these extreme cases are well-defined, there exists a vast diversity of topologies between these fully-connected extremes. Graphs can be: random, where any two nodes are connected following an established probability, leading to a Poisson degree distribution [44]; small-world, where the topology is random with internal cycles that connect otherwise far apart nodes; scale-free, where a handful of nodes have exponentially large number of neighbors, while the majority of nodes only have a few, leading to an exponential degree distribution [44]; regular, where all nodes have the same or nearly the same number of neighbors; and multipartite, with 2 or more sets of nodes that are not connected within their set, but exclusively between sets.
@@ 4.2 Generated Tasks
We also articulate new tasks that have not appeared in previous network task taxonomies, showing the generative power of our taxonomy. For example, “How many diameter endpoint nodes does the graph have?” (T17), arises from the Constituent scope, a Quantify action, and the diameter endpoint node target. The concrete task of “Where in the graph are the bridge edges?” (T14) also occurs at the Constituent level, with a Locate action on the bridge edge target
@@ 4.4 Ecological Validity of Tasks
5 Related Work
We discuss previous work on general task taxonomies, network task taxonomies that pertain to the Core level of our data abstraction, and network task taxonomies that cover semantically richer data abstractions.
5.1 General Task Taxonomies
An early taxonomy from Shneiderman [59] provides a high-level breakdown of tasks and of data types, while Amar et al [2] provide a low-level taxonomy of analytic tasks. In 2013 Brehmer and Munzner [12] proposed bridging from low-level tasks to higher-level tasks with a multi-level typology of abstract tasks. Munzner [42] breaks down tasks into action-target pairs; we extend this approach in our taxonomy.
From this datatype-agnostic foundation, many other taxonomies focus on narrower, more specialized scopes of specific data types. For example, Valiati et al [64] cover multidimensional data. Nusrat and Kobourov cover the intersection of geographic data and the visual encoding technique of cartograms [46]. None of these papers focus on the data type of networks, which is the subject of our work.
5.2 Network Task Taxonomies
The seminal network task taxonomy from Lee et al [35] includes some topology-based tasks that are well-suited for unlabelled graphs, which we incorporate into our list of applicable concrete tasks. However, the majority of their topology tasks begin with a “given a set of nodes” as the starting point, which requires meaningful semantic labels that are not available with unlabelled graphs. Moreover, they do not cover the Multiple or Pair scopes at all.
Although the taxonomy by Kerracher et al [30] primarily focuses on temporal graphs, they also discuss the more general issue that structural tasks with no attribute metadata involved are an important category, many of which we include in our list. They identify comparison as a task of interest, corresponding to our Compare action at the Pair scope. However, they do not address the Multiple scope.
Some structural comparison tasks relevant to unlabelled graphs are described by Bae et al [4], but the main focus of their paper is the perceptual mechanisms involved in network pattern detection.
While some aspects of these taxonomies cover tasks that pertain to the Core data abstraction layer (Fig. 2-A), their main focus is tasks that primarily rely on semantic data at higher levels of semantic richness, and none have comprehensive coverage of the full set of tasks that apply to unlabelled graphs.
5.3 Labelled Graph Task Taxonomies
Many specialized task taxonomies address the types of graphs that we assign to the formalism layer of our data abstraction, including dynamic graphs that focus on changes over time of attribute-enriched networks [1, 9, 30], heterogeneous graphs with diverse multivariate attributes [52, 45], and compound graphs with group semantics as node attributes [57, 65]. Others focus on domain-specific analysis, such as biological pathway networks [43], with tasks that rely on rich semantic attributes. However, none of these tasks are applicable to unlabelled graphs. A comprehensive 2023 survey of network visualization task taxonomies [20] illustrates how much of the previous work is dedicated to networks with attributes, confirming the need for a task taxonomy for the unlabelled graph context.
We can also explore how real-world constraints shape the kinds of tasks researchers prioritize in practice. A practitioner survey by Sahu et al [56] highlights that large-scale network analysis often involves graphs exceeding a billion edges, with dynamic structures and complex label attributes.
Figure 4: The 6 visual encodings that we consider for our assessment of unlabelled graph tasks. All of these example views share the same underlying graph topology: ’moreno-health’ from the KONECT repository [34], an empirical social network of 2539 nodes (people) and 10455 edges (friendships).
6 Analysis: Visualization Assessment
We validate the evaluative power of our taxonomy by using it to guide our assessments about the effort and success of a selection of 17 unlabelled graph tasks applied to 6 visual encodings. We explain the chosen encodings, elaborate on the analysis process, and summarize the analysis results.
@@ 6.1 Visual Encodings
Adjacency Matrix (AM). In the AM visual idiom each node is assigned a unique row and column in a matrix, where non-zero cells at the intersection of two connected nodes represent their shared edge. Among the visual idioms we consider, AM is the only lossless one: the entire graph topology can be reconstructed from the matrix data table. However, salient visual patterns still depend on the seriation, or permutation method, of the rows and columns; in Fig. 4 we use generic hierarchical clustering. Revealing network structure through matrix seriation is a non-trivial process that leads to varied outcomes [10, 61].
@@
Node-Link (NL). We also analyze a basic version of node-link layout that does not use any information beyond the core topology, with straight edge lines and equal-sized node points, monochromatic, and spatially arranged using a generic force-directed layout [32]. Although many more sophisticated node-link approaches have been proposed, most heuristics rely on attribute values, which are unavailable in the unlabelled graph context.
Hop-Census Plot (HC). In previous work, we [47] introduced and evaluated Census vectors as an invariant descriptor data structure that effectively captures topological information. We proposed HC plots as an idiom where each Census vector is visually encoded as a polyline, and the collection of polylines form a variant of a parallel coordinate plot where the order of the axes is dictated by the data structure itself. The HC we consider in this paper is based on Census-Stub vectors, each computed from generalized stub (half-edge) degrees.
Census-Census Plot (CC). We [47] also noted that Census vectors can be independently computed for generalized node degree (Census-Node vectors) and generalized stub degree (Census-Stub vectors). We proposed CC plots to visually encode differential trajectories from respective Census vectors. In this plot we encode Census-Node orthogonal to Census-Stub.
@@
BMatrix Network Portrait (NP). The NP idiom proposed by Bagrow et al [6, 5] uses a matrix heatmap to encode aggregated information about unlabelled graphs, built upon a data structure of polylines equivalent to the Census-Node vectors we discuss above. In contrast to the HC plots which directly draw polylines, this idiom aggregates the information they contain into discretized colored cells that can reveal regions of density.
@@
Graph Thumbnail (GT). The GT idiom [70] uses concentric bubbles to display information arising from a k-core decomposition algorithm, aggregating nodes of the same k-core level into a single bubble, and color encoding levels with a heatmap from outermost peripheral 1-core (blue), to inner 2-core (green), to 3-core (bright orange), and so on to the innermost cores (dark orange). The nested structure of GT arises from a trade-off of information expressiveness that maximizes the salience of a specific signal to highlight core-periphery structure, while letting go of information fidelity: the original topology is aggressively aggregated and cannot be reconstructed from GT.
6.2 Process: Analysis
@@
For each possible combination of abstract Scope + Action enumerated in the taxonomy (Sec. 3), we sampled one or two Target possibilities to select one or two concrete tasks. In some cases our choice was guided by tasks that are particularly well handled by one of the invariant encodings, to illuminate the interesting possibilities in these under-studied approaches. We then assessed the performance of each visual encoding according to two criteria: the Effort required from the viewer to carry out the task, and the likelihood of Success across a range of datasets with different characteristics.
We did this assessment for two sizes of graph data: Small-scale (<100 nodes) and Large-scale (>1000 nodes), a size distinction following survey results [16]. Our goal was to roughly divide visual encodings according to contexts where each works well, against scenarios where they begin to fail. For example, large graphs would have so much clutter when using basic force-directed placement that they would appear to be hairballs [42]. We consider relatively sparse graphs that are common in visualization contexts, where there are no more than three times the number of edges vs nodes [41].
To avoid a coarse binarized pass or fail assessment, we iterated and settled into 4 bins so our results capture more nuance of our consensus on visual encoding performance. For Effort, our bins are Easy, Middling, Difficult, and Impossible; and for Success our bins are Always, Sometimes, Rarely, and Never. Our qualitative assessment is both preliminary and subjective, and should be treated as a starting point rather than a definitive statement of facts: our intent is to spur future work that uses the taxonomy to conduct empirical studies.
@@
The first author carried out the first round of this assessment through inspection of layouts for each of the encodings across an eclectic collection of 81 unlabelled graphs. This benchmark dataset was already demonstrated on five of the encodings in our own previous work [47]; we added Graph Thumbnails. The supplemental material contains collages of these 81 layouts for all six encodings. All authors discussed the assessment, handling disagreements through iterative rounds of analysis until convergence.
Figure 5: We illustrate the evaluative power of our taxonomy by selecting 17 unlabelled graph tasks covering all possible Scope + Action combinations with one or two sampled from Target (rows) to consider with respect to 6 network visualizations (columns). At each task and visual encoding intersection, the table shows our assessment of user performance vs. data scale with a quadrant heatmap: top left for task effort on Small graphs, top right for task success on Small graphs, bottom left for task effort on Large graphs, and bottom right for task success on Large graphs. Task effort and success are indicated with a diverging blue-red colormap mapped to 4 ranked assessment levels: Easy, Middling, Difficult, and Impossible (for task effort); and Always, Sometimes, Rarely, and Never (for task success).
@@ 6.3 Analysis Results
Fig. 5 shows the summarized analysis results; a full spreadsheet is available as supplemental materials with a brief rationale for each cell.
@@
The Adjacency Matrix idiom (first column) has low performance for the multi-graph scopes because of its reliance on matrix seriation. In the unlabelled graph context, where there are no persistent labels, there is no way to find correspondences between nodes across graphs, so matrix seriation is unlikely to reveal similarities between them that could be revealed through visual patterns (T01-T05). The Adjacency Matrix idiom excels at a few tasks at the Constituent level, namely the Quantify (T16) and Classify (T15) tasks for dead-end node targets, where counting and finding matrix rows that have only one cell is easy for small graphs and middling for large graphs. The well-known challenge of Adjacency Matrix at handling tasks that involve tracing topological paths is apparent for many of the narrow-scope targets, including bridges (T14), chains (T12), and diameters (T09, T10, T17).
@@
The Hop-Census Plot and Census-Census Plot invariant descriptor idioms (third and fourth columns) also do best at the multi-graph Multiple and Pair broad-scopes, in part due to their shared frame of reference affordance [47]. They also have some resilience to scale changes in those scopes. The Hop-Census idiom has good performance at tasks involving diameter targets at the Single scope (T09, T10) as does the BMatrix Network Portrait, as these idioms directly visually encode the relevant information in a way that can be easily counted for Quantify actions and spotted for Classify actions. Hop-Census is unique among the invariant idioms in also performing well at many tasks in the Subgraph and Constituent scopes, including diameter (T17) and dead-end (T13, T15, T16) targets. It also performs reasonably well for the chain target at the Subgraph scope, which can sometimes be spotted with middling effort because topological chains have a distinct constant-value polyline across parallel hop axes. The same performance applies to Node-Link views for this task, where chains may be visible in the layout.
@@
For “Does this graph have scale-free structure” (T07), the Adjacency Matrix view can sometimes handle small graphs with middling effort using the appropriate matrix seriation, where the scale-free pattern can be made salient by sorting according to number of neighbors and looking for an exponential decay pattern; with large graphs, the effort is large and the success is rare as the matrix grows large and the pixels grow small. Notably, the challenges in seriation at the broad-scope does not apply here, because there is no need to find node correspondences between graphs. The Node-Link view would require middling effort for small graphs, where the scale-free pattern could be salient in rare cases of forgiving topology, and would not work at all in the large cases. The Census-Census plots can sometimes verify scale-free structure for small graphs with middling effort, although they more easily rule out when the topology is not scale-free; for large graphs, occlusion leads to difficult status, although sometimes scale-free patterns can still surface. The aggregate structure of the BMatrix Network Portrait and Graph Thumbnail views leads to difficult effort and rare success at both scales.
@@
Our assessment summary (Fig. 5) also shows that the effect of Action on idiom performance is often small, for tasks which share both Scope and Target. In some concrete cases, the Classify task is easier than the Quantify task, as with noting the existence of a dead-end node target (T15) vs counting the number of instances of it within the graph (T16).

### v2

@@ 5.2 Action
Locate. Asks “where are the targets situated within the graph?”, and applies them to the Subgraph and Constituent scopes. The location may involve proximity between targets of the same type, or between two different kinds of targets.
@@ 5.3 Targets
At the Multiple and Pair scopes, the target is simply the graph itself; that is, the Unlabelled graph topology (Sec. 3) of a simple graph consisting of a single connected component. At the Pair scope there are two such targets, and at the Multiple scope there are many.
At the Single taxonomy scope, the target pertains to the entire graph, encompassing all of its nodes and edges. Many of the targets are names from the graph theory literature that describe global topological characteristics of a graph. A graph of $N$ nodes is a tree when it has N1N-1 edges and a complete graph when it has $N\choose 2$} edges. While these extreme cases are well-defined, there exists a vast diversity of topologies between these fully-connected extremes. Graphs can be: random, where any two nodes are connected following an established probability, leading to a Poisson degree distribution [6]; small-world, where the topology is random with internal cycles that connect otherwise far apart nodes; scale-free, where a handful of nodes have exponentially large number of neighbors, while the majority of nodes only have a few, leading to an exponential degree distribution [6]; regular, where all nodes have the same or nearly the same number of neighbors; and multipartite, with 2 or more sets of nodes that are not connected within their set, but exclusively between sets.
@@ 6.2 Generated Tasks
We also articulate new tasks that have not appeared in previous task taxonomies, showing the generative power of our taxonomy. For example, “How many diameter endpoint nodes does the graph have?” (T17), arises from the Constituent scope, a Quantify action, and the diameter endpoint node target. The concrete task of “Where in the graph are the bridge edges?” (T14) also occurs at the Constituent level, with a Locate action on the bridge edge target.
@@ 6.4 Ecological Validity of Tasks
7 Example Analysis
We validate the descriptive and evaluative power of our taxonomy by using it to guide an example analysis in which we make educated guesses about the effort and success of a selection of 17 unlabelled graph tasks applied to 6 visual encodings. Our intent is to demonstrate that the taxonomy has good coverage of the breadth of tasks that pertain to unlabelled graphs, and enough precision to support the analysis of visual encodings.
We explain the chosen encodings, elaborate on the analysis process, and summarize the analysis results.
@@ 7.1 Visual Encodings
Figure 4: The six visual encodings we consider for our assessment of unlabelled graph tasks. These six views share the same underlying graph topology: ’moreno-health’ from the KONECT repository [73], an empirical social network of 2539 nodes (people) and 10455 edges (friendships).
@@
Adjacency Matrix. Among the visual idioms we consider, Adjacency Matrix is the one with the highest information fidelity: the entire graph topology can be reconstructed from the visual representation of the matrix data table. In the Adjacency Matrix visual idiom each node is assigned a unique row and column in a matrix, where non-zero cells at the intersection of two connected nodes represent their shared edge. This idiom does not require any information above the Unlabelled layer. When the represented graph topology is unlabelled the graph is necessarily undirected, so the visual patterns are always diagonally symmetric because there is no distinction between edge A-to-B and edge B-to-A. However, despite diagonal symmetry, surfacing salient visual patterns still depends on matrix seriation, namely the row and column permutation method, which is an algorithmic design decision. In Fig. 4 we use generic Louvain communities to determine matrix seriation [74]. Revealing network structure through the visual patterns that arise from matrix seriation is a non-trivial process that leads to varied outcomes [24, 40].
@@
Node-Link. We also analyze a basic version of Node-Link layout that again does not use any information beyond the Unlabelled graph topology, with edges represented by straight-line segments and nodes as equal-sized disks, monochromatic, and spatially arranged using a generic force-directed layout [22]. Although many more sophisticated node-link approaches have been proposed, most heuristics rely on Attributed values, which are unavailable in the Unlabelled context (Fig. 2). Like matrix seriation in Adjacency Matrix, the spatial embedding of the nodes in a node-link diagram is also a non-trivial process, often with extremely diverse visual outcomes.
Hop-Census Plot. In our previous work [3], we proposed a new line-based invariant plot we call Hop-Census, which encodes the Census data structure, which is composed of one vector of integers per node, and each vector is visually encoded as a polyline in a shared coordinate system. The vector integers represent counts of elementary constituents found during a traversal of the graph from a source node. The horizontal axis is bounded by the diameter of the graph. The vertical axis shows the counts of the elementary constituents, for example stub (half-edge) degree in the instantiation shown in Fig. 4. The polyline for a given node captures the number of constituents at each hop: that is, the number of half-edges adjacent to the source node itself at the first hop, then the number of half-edges adjacent to the neighboring nodes at the second hop, and so on. Therefore, the Hop-Census plot is a type of parallel coordinate plot, where the horizontal ordering of the vertical axes is dictated by the data itself rather than being arbitrarily determined. Moreover, the extent of each vertical axis is uniform, providing an absolute coordinate system in which to visually embed the Census vector polylines. Thus, this plot shows a collective multi-polyline visual pattern that is itself invariant (i.e. always the same for all isomorphs of a graph).
Census-Census Plot. In our previous work [3], we also proposed Census-Census plots as another example of a line-based invariant plot. Census vectors can count different kinds of elementary constituents: stubs, nodes, and edges. These plots show two of these different kinds of Census vectors orthogonal to each other. The path traced by a single polyline follows a trajectory that reveals the differences between these different constituent counts, as the graph is traversed. In the instantiation shown in Fig. 4, the horizontal axis is node degree and the vertical axis is stub degree. The length of a Census vector is always the same no matter which kind of constituent is counted, leading to an absolute shared coordinate system that affords this orthogonal juxtaposition. Like Hop-Census plots, the collective multi-trajectory visual pattern is invariant.
@@
BMatrix Network Portrait. The Network Portrait idiom proposed by Bagrow et al. [4, 16] is another invariant plot that uses a matrix heatmap to encode aggregated information about Unlabelled graphs, built upon a data structure of polylines equivalent to the Census vectors we discuss above. In contrast to the Hop-Census plots which directly draw polylines, the BMatrix Network Portrait heatmap idiom aggregates the information they contain into discretized colored cells that can reveal regions of density. The axes of the Network Portrait are transposed from the Census-based plots: the horizontal axis is generalized node degree, and the vertical axis is hop distances.
@@
Graph Thumbnail. While Census plots and the Network Portrait rely on hop-based invariant data structures (Census vectors and BMatrix, respectively), the Graph Thumbnail idiom [5] relies on a different invariant data structure: the k-core decomposition of the graph. The definition of a k-core subgraph is that the value k is equal to the minimum degree of all nodes within it, which effectively decomposes the graph into nested subgraphs; every node within a k-core subgraph has at least degree k. Graph Thumbnails uses concentric bubbles to aggregate nodes of the same k-core subgraph into a single bubble, and then color encodes k-core hierarchy with a heatmap from outermost peripheral 1-core (blue), to inner 2-core (green), to 3-core (bright orange), and so on to the innermost cores (increasingly darker orange hues). The nested structure of Graph Thumbnails arises from a trade-off of information expressiveness that maximizes the salience of the graph’s k-core decomposition to highlight core-periphery structure, while letting go of information fidelity through the choice of concentric bubble visual encoding: the original topology is aggregated and impossible to reconstruct from the visual idiom.
7.2 Analysis Dimensions
For each possible combination of abstract Scope + Action enumerated in the taxonomy (Sec. 5), we sampled one or two Target possibilities to select one or two concrete tasks. The resulting list of tasks provides evidence of the descriptive power of the taxonomy in terms of coverage. We see that the Scope dimension leads to good coverage across ranges of analysis granularity; as we discuss in Sec. 4, previous taxonomies do not provide adequate coverage of the Pair and Multiple levels. This triplet-based analysis provides broad coverage of the full suite of unlabelled graph considerations; previous task taxonomies that focus on labelled graph tasks do not provide such a way to systematically enumerate unlabelled tasks.
@@ 7.2 Analysis Dimensions
In some cases our choice of which task to sample for the triplet was guided by tasks that are particularly well handled by one of the invariant plots, to illuminate the interesting possibilities in these under-studied invariant techniques. We then made educated guesses about the performance of each visual encoding according to two criteria: the Effort required from the viewer to carry out the task, and the likelihood of Success across a range of datasets with different characteristics.
@@
The first author carried out the first round of assessment through inspection of each of the six visual encodings across an eclectic collection of 81 unlabelled graphs. This benchmark dataset was already demonstrated on five of the encodings in our own previous work [3]; we added Graph Thumbnails. Our Supplemental materials contain the six 9-by-9 collages that showcase the 81 networks from our eclectic benchmark at once; one collage per each of the six visual encoding. All authors discussed the assessment, handling disagreements through iterative rounds of analysis until convergence.
To avoid a coarse binarized pass-or-fail assessment, we iterated and settled into 4 bins so our results capture more nuance of our educated guess consensus on visual encoding performance. Our analysis dimensions Effort and Success helped us better capture our different points of view simultaneously and arrive at consensus. For the Effort dimension (i.e. our estimate of the cognitive load demanded from human viewers) our bins are Easy, Middling, Difficult, and Impossible. For the Success dimension (i.e. our estimate of the likelihood of task completion) our bins are Always, Sometimes, Rarely, and Never. This qualitative assessment is both preliminary and subjective, and should be treated as a starting point rather than a definitive statement of facts: our intent is to spur future work that uses the taxonomy to conduct empirical studies.
We also consider a Scale dimension (i.e. the cardinality of the graph) with two sizes of graph data: Small-scale (<100 nodes) and Large-scale (>1000 nodes), a size distinction following survey results [38]. Our goal was to roughly divide visual encodings according to contexts where each works well, against scenarios where they begin to fail. For example, large graphs would have so much clutter when using basic force-directed placement that they would appear to be hairballs [9]. We consider relatively sparse graphs that are common in visualization contexts, where there are no more than three times the number of edges vs nodes [19].
@@ 7.3 Analysis Results
Fig. 5 shows the summarized analysis results; a full spreadsheet is available in Supplemental materials with a brief rationale for each cell.
Figure 5: We illustrate the descriptive and evaluative power of our taxonomy by selecting 17 unlabelled graph tasks covering all possible Scope + Action combinations with one or two sampled from Target (rows) to consider with respect to 6 network visualizations (columns). At each task and visual encoding intersection, the table shows our assessment of user performance vs. data scale with a quadrant heatmap: top left for task effort on Small graphs, top right for task success on Small graphs, bottom left for task effort on Large graphs, and bottom right for task success on Large graphs. Task effort and success are indicated with a diverging blue-red colormap mapped to 4 ranked assessment levels: Easy, Middling, Difficult, and Impossible (for task effort); and Always, Sometimes, Rarely, and Never (for task success).
@@
We consider that the Adjacency Matrix idiom (first column) has a low performance for the multi-graph scopes for two reasons: first, we expect users to have perceptual and cognitive difficulties reading the Adjacency Matrix visual pattern [75]; and second, because the visual pattern itself relies on a non-trivial matrix seriation choice [40]. In the unlabelled graph context, where there are no persistent labels, there is no way to find correspondences between nodes across graphs, so matrix seriation is unlikely to reveal similarities between them that could be revealed through visual patterns (T01-T05). The Adjacency Matrix idiom excels at a few tasks at the Constituent level, namely the Quantify (T16) and Classify (T15) tasks for dead-end node targets, where counting and finding matrix rows that have only one cell is easy for small graphs and middling for large graphs. The well-known challenge of Adjacency Matrix at handling tasks that involve tracing topological paths is apparent for many of the narrow-scope targets, including bridges (T14), chains (T12), and diameters (T09, T10, T17).
@@
The Hop-Census Plot and Census-Census Plot invariant plot idioms (third and fourth columns) also do best at the multi-graph Multiple and Pair broad-scopes, in part due to their shared frame of reference affordance [3]. They also have some resilience to scale changes in those scopes. The Hop-Census idiom has good performance at tasks involving diameter targets at the Single scope (T09, T10) as does the BMatrix Network Portrait, as these idioms directly visually encode the relevant information in a way that can be easily counted for Quantify actions and spotted for Classify actions. Hop-Census is unique among the invariant idioms in also performing well at many tasks in the Subgraph and Constituent scopes, including diameter (T17) and dead-end (T13, T15, T16) targets. It also performs reasonably well for the chain target at the Subgraph scope, which can sometimes be spotted with middling effort because topological chains have a distinct constant-value polyline across parallel hop axes. The same performance applies to Node-Link views for this task, where chains may be visible in the layout.
@@
For “Does this graph have scale-free structure” (T07), the Adjacency Matrix view can sometimes handle small graphs with middling effort using the appropriate matrix seriation, where the scale-free pattern can be made salient by sorting according to number of neighbors and looking for an exponential decay pattern; with large graphs, the effort is large and the success is rare as the matrix grows large and the pixels grow small. Notably, the challenges in seriation at the broad-scope does not apply here, because there is no need to find node correspondences between graphs. The Node-Link view would require middling effort for small graphs, where the scale-free pattern is more likely to be salient, but would not work at all in the large cases. The Census-Census plots can sometimes verify scale-free structure for small graphs with middling effort, although they more easily rule out when the topology is not scale-free; for large graphs, occlusion leads to difficult status, although sometimes scale-free patterns can still surface. The aggregate structure of the BMatrix Network Portrait and Graph Thumbnail views leads to difficult effort and rare success at both scales.
@@
Our assessment summary (Fig. 5) also shows that the effect of Action on idiom performance is often small, for tasks which share both Scope and Target. In some concrete cases, the Classify task is easier than the Quantify task, as with noting the existence of a dead-end node target (T15) versus counting the number of instances of it within the graph (T16).